What is the ratio of the –COOH group to the –COO-1 group for proline at pH 7.0?

Calculate the pI of A) Tyrosine and B) Arginine.

What is the overall charge of the pentapeptide SHDNF at pH 7.0?

What is the overall charge of the pentapeptide SHDNF at pH 3.0 (assume no hydrolysis of the backbone occurs)?

What is the overall charge of the pentapeptide SHDNF at pH 10.0 (assume no hydrolysis of the backbone occurs)?
Answer #1

1. We wish to determine the ratio of COOH and COO- groups in proline at pH 7.0.

The pKa of the acid (-COOH) group in proline is 1.99.

Use the Henderson-Hasslebach equation to determine the ratio of COO- and COOH in proline at pH 7.0.

pH = pKa + log [COO-]/[COOH]

=====> 7.0 = 1.99 + log [COO-]/[COOH]

=====> 5.01 = log [COO-]/[COOH]

=====> [COO-]/[COOH] = antilog (5.01)

=====> [COO-]/[COOH] = 1.023*105

The ratio of COO- to COOH in proline at pH 7.0 is 1.023*105:1; i.e, the carboxyl group is proline is completely deprotonated to the –COO- form.

However, the question asks for the ratio of –COOH to –COO- at pH 7.0 which is simply the reverse of the above ratio, i.e.,

[COOH]/[COO-] = 1/(1.023*105/1)

= 1/(1.023*105)

= 1:9.775*10-6

≈ 1:9.8*10-6 (ans).